# LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation

### v1

@@ Abstract
Speculative decoding (SD), where a small draft model is employed to propose draft tokens in advance and then the target model validates them in parallel, has emerged as a promising technique for LLM inference acceleration. Many endeavors to improve SD are to eliminate the need for a draft model and generate draft tokens in a retrieval-based manner in order to further alleviate the drafting overhead and significantly reduce the difficulty in deployment and applications. However, retrieval-based SD relies on a matching paradigm to retrieval the most relevant reference as the draft tokens, where these methods often fail to find matched and accurate draft tokens. To address this challenge, we propose LogitSpec to effectively expand the retrieval range and find the most relevant reference as drafts. Our LogitSpec is motivated by the observation that the logit of the last token can not only predict the next token, but also speculate the next next token. Specifically, LogitSpec generates draft tokens in two steps: (1) utilizing the last logit to speculate the next next token; (2) retrieving relevant reference for both the next token and the next next token. LogitSpec is training-free and plug-and-play, which can be easily integrated into existing LLM inference frameworks. Extensive experiments on a wide range of text generation benchmarks demonstrate that LogitSpec can achieve up to 2.61$\times$ speedup and 3.28 mean accepted tokens per decoding step. Our code is available at https://github.com/smart-lty/LogitSpec.
@@ 1 Introduction
To address this challenge, speculative decoding (SD) (Leviathan et al., 2023; Chen et al., 2023) has emerged as a promising approach for lossless LLM inference acceleration. The key idea of SD is to employ a small draft model to first generate $\gamma$ draft tokens, and then the target model validates these tokens in parallel within a single forward, ensuring the output distribution to be unchanged. However, deploying an extra draft model introduces considerable overhead and difficulties, including (a) complex implementation and deployment, especially in integration with other techniques,(Liu et al., 2025) (b) additional memory overhead, especially in long-context scenarios. Moreover, when there exists no available draft model for SD, it takes substantial training cost to construct a compact draft model (Li et al., 2025a). Therefore, many existing works dedicate to developing draft-model-free SD methods (Saxena, 2023; Fu et al., 2024; He et al., 2024), where the draft tokens are generated in a retrieval-based manner. At each decoding step, they retrieve the relevant reference according to the last few tokens and extract draft tokens from the reference. In this way, they can explicitly eliminate the need for a draft model and reduce the drafting overhead.
Figure 1: Illustration of vanilla retrieval-based SD method and our LogitSpec. (a) Retrieval-based SD retrievals the wrong token “triangle” according to the next token “the”. (b) LogitSpec first speculates the next next token “area”, and then retrievals the right relevant reference “area of” accord to “the area”. This simple example illustrates how LogitSpec utilizes the last logit to speculate the next next token and improves the retrieval accuracy.
@@
In this paper, we propose a simple yet effective retrieval-based SD framework, namely LogitSpec, which leverages the logit of the last token (last logit) to predict the next next token and improves the accuracy of the retrieval reference. LogitSpec is motivated by the observation that the last logit can speculate the next next token with a relatively high accuracy. Based on this observation, we use the speculated next next token as a guidance to retrieval reference. Specifically, LogitSpec generates draft tokens in two steps: (1) utilizing the last logit to speculate the next next token; (2) retrieving relevant reference for both the next token and the next next token. As shown in Figure 1 (b), with the speculated token “area”, LogitSpec successfully retrievals the correct draft tokens “area of”. This two-step process enables better retrieval accuracy—it can help filter the most relevant reference when the reference is redundant, while extending the searching space when there exists no relevant reference. To summarize, our contributions are:
@@
    We empirically observe that the last logit can speculate the next next token with a relatively high accuracy. Unlike other retrieval techniques which are sensitive to the specific task, this property is robust and effective across various tasks.
@@
    We conduct extensive experiments on various text generation benchmarks to demonstrate the effectiveness of LogitSpec. Notably, LogitSpec achieves up to 2.61$\times$ speedup and 3.28 mean accepted tokens per decoding step without the need for an extra draft model.
@@ 2 Related Work
This line of research focuses on employing an extra draft model for speculative decoding. Some existing works make their efforts on developing a more compact draft model with less drafting overhead. Medusa (Cai et al., 2024) pioneers this direction by adding extra MLP heads at the top of the target model, reusing the hidden states from the target model and predicting the next few tokens in parallel. However, the generation from these decoding heads is independent, which harms the accuracy of the draft tokens. Therefore, some works (Ankner et al., 2024; Xiao et al., 2024) propose to add sequential dependencies for better performance. Glide (Du et al., 2024) takes the similar idea of Medusa and propose to reuse KV cache from the target model. Recently, Eagle (Li et al., 2025b, 2024, a) has dominated this research line by training an auto-regressive head and scaling up the training data. However, while these methods achieve superior speedup, all of them rely on an extra draft model, which necessitate extra parameters or extensive training. The existence of the draft model significantly limits their application: (a) deploying a draft model requires complex and careful implementation. (b) Both the draft model and its KV cache require additional GPU memory, which may incur load unbalance. In the long context settings, even the KV cache for the draft model may exceed the capacity of 1 GPU and leads to significant resource competitions; (c) the training cost is substantial, and the draft model requires retraining when the target model is updated.
@@
This line of research focuses on eliminating the need for a draft model. Some existing works (Zhang et al., 2024a; Elhoushi et al., 2024; Xia et al., 2025) observe the layer sparsity of LLMs that not all the layers are necessary to predict the next token. They propose methods to generate draft tokens with a subset of the layers of target model. However, the number of skipped layers is limited, which affects the overall speedup. Recently, retrieval-based SD (He et al., 2024; Saxena, 2023; Fu et al., 2024) has dominated the draft-model-free speculative decoding by retrieving the relevant reference as the draft tokens. The relevant reference can be derived from the question and generated tokens, or external database. However, there exists two common issues during the retrieval process: (a) retrieval from the context often fails to find matched draft tokens; (b) retrieval from external database often struggles to find accurate draft tokens. These two issues motivate us to develop a more appropriate retrieval-based speculative decoding framework.
@@ 3 Background
Let $\bm{x}$ denote an input sequence (prefix), a speculative decoding step consists of a drafting phase and a verification phase. During the drafting phase, the draft model $\mathcal{M}_{q}$ is employed to give $\gamma$ draft tokens $x_{1},x_{2},...,x_{\gamma}$ by running $\gamma$ times model forward and sample. Here, we denote the output logit $\mathcal{M}_{q}(\bm{x}+[x_{1},...,x_{i-1}])$ as $q_{i-1}$, then each draft token is given by $x_{i}\sim q_{i-1},i=1,...,\gamma$. During the verification phase, the prefix $\bm{x}$ together with $\gamma$ draft tokens are sent to $\mathcal{M}_{p}$ for verification. The target model $\mathcal{M}_{p}$ inputs $\bm{x}+[x_{1},...,x_{\gamma}]$ and outputs the logits $p_{0},p_{1},...,p_{\gamma}$. Then SD sequentially verifies $x_{i}$ via speculative sampling (Leviathan et al., 2023; Chen et al., 2023), where the acceptance rate is given by:
$\alpha_{i}=\left\{\begin{aligned} &1\quad\qquad\quad p_{i-1}[x_{i}]\geq q_{i-1}[x_{i}],\\ &\frac{p_{i-1}[x_{i}]}{q_{i-1}[x_{i}]}\quad p_{i-1}[x_{i}]<q_{i-1}[x_{i}],\\ \end{aligned}\right.$ (1)
@@
=$\mathcal{R}=\{(x_{i}^{1},x_{i}^{2},...,x_{i}^{n})\}_{i=1}^{N}.$ (2)
Here, $N$ denotes the total number of $n$-grams in $\mathcal{R}$. Existing retrieval-based SD methods mainly differ in constructing $\mathcal{R}$. At each SD decoding step, given a query $m$-gram tokens $x_{q}^{1},x_{q}^{2},...,x_{q}^{m}$, the retrieval model first traversals the reference, finds matched $n$-grams and returns the subsequent tokens of the matched tokens:
$\textsc{Match}(\mathcal{R},(x_{q}^{1},x_{q}^{2},...,x_{q}^{m}))=\{(x_{i}^{m+1},...,x_{i}^{n})\}_{i\in\mathcal{S}},$ (3)
@@ 4 Method
Figure 2: Motivated observations. (a) The last logit can speculate the next next token with a relatively high accuracy. In over 50% decoding steps, the next next token can be found in the top-60 entries within the last logit. This prediction ability exists consistently across different model sizes and architectures. (b) Compared with other retrieval-based methods, the prediction of last logit demonstrates superior robustness to the downstream task type. Both of these observations motivate us to utilize the last logit to guide the retrieval process.
@@ 4.1 Motivated Observation
Drawing insights from multi-token prediction (MTP) (Gloeckle et al., 2024) that fine-tuned LLMs can predict multiple tokens in a single forward, we are interested in the ability of LLMs to predict the next next token without fine-tuning. We conduct a simple experiment to investigate the rank of the next next token in last logit, which is used to predict the next token. Specifically, we use a small fraction of Spec-Bench (Xia et al., 2024) that contains 13 sub-tasks (detailed in Appendix B) and 6 different LLMs as the backbone models. The experimental results, as shown in Figure 2, demonstrate that the last logit has a strong potential to predict the next next token. In over 50% decoding steps, the next next token can be found within the top-60 entries of the last logit. For Llama 3.1 8B (Team, 2024) with a large 128K vocabulary, this percentage even increases to 64% within top-60 entries of the last logit. The results of 6 different LLMs on 13 downstream tasks demonstrate that this observation holds consistently across different models (Vicuna, Llama 3.1 and Qwen) and scales (ranging from 0.5B to 33B).
We further conduct experiments on Spec-Bench with Vicuna 7B to investigate the robustness of the predictive capability of the last logit across different tasks. We propose a minimal speculative decoding algorithm with last logit, last logit decoding for this experiment, where we directly use the top-60 entries of the last logit to serve as the draft tokens. Each draft token is a guess to the next next token. If the target model accept any one of the draft tokens, the others are dropped. Consequently, at each decoding step, the accepted length is either 2 or 1. As shown in 2, this simple SD method yields mean accepted tokens per decoding step (MAT) of over 1.5. Compared with other retrieval-based SD methods such as PLD (Saxena, 2023) and REST (He et al., 2024), which are highly affected by the task type, the MAT of last logit decoding exhibits superior robustness to the task. These results demonstrate that the ability of the last logit to predict the next next token widely exists in modern LLMs themselves, and strongly motivate us to improve retrieval-based SD methods with last logit. We provide a detailed experimental setup for the motivation experiments in Appendix B.
Figure 3: An overview of LogitSpec. At each decoding step, LogitSpec first utilizes the top-$k$ entries of the last logit as the speculation to the next next token. Then, LogitSpec retrievals relevant references for both the next token and the next next token. Finally, LogitSpec organizes the draft tokens into a draft tree and prepares a tree attention for parallel verification.
@@ 4.2 LogitSpec Drafting
[…]
=$\mathcal{L}=\{\widetilde{x}_{i+2}\ |\ \widetilde{x}_{i+2}\in\textsc{Top}_{k}(p_{i})\land\widetilde{x}_{i+2}\neq x_{i+1}\}.$ (4)
Then, we use the retrieval model $\mathcal{R}$ to retrieval reference for both the next token and next next tokens:
$\mathcal{D}=\textsc{Match}(\mathcal{R},\overbrace{(x_{i-m+2},...,x_{i},x_{i+1})}^{m\text{-gram for next token}})\cup\{\textsc{Match}(\mathcal{R},\overbrace{(x_{i-m+3},...,x_{i},x_{i+1},\widetilde{x}_{i+2})}^{m\text{-gram for next nexttoken}})\}_{\widetilde{x}_{i+2}\in\mathcal{L}}$ (5)
Here, $m$-gram denotes the last $m$ tokens as the query to retrieval the matched reference. In our experiments, we retrieval draft tokens with $m=3$ first. If no matched tokens are found, we decrease $m$ to 2 and so on. For simplicity, we only use the user-input prompt and decoded tokens to construct the retrieval model $\mathcal{R}$, that is to say, our LogitSpec is query-independent, and will not be affected by other queries. We give the detailed implementation of retrieval process in Appendix D.1.
Taking the example in Figure 3, we first select the top-$k$ entries of the last logit as the speculation to the next next token. Then, we retrieval reference for both the next token “the” and the next next token “area”, and return “of the”. After that, we organize these draft tokens into a draft tree, where the next token serves as the root. Finally, token sequence “area of the” is accepted by $\mathcal{M}$.
After the retrieving process, we employ a simple pruning strategy to control the number of the draft tokens. Specifically, we do not prune the retrieval tokens for the next token. For each speculated next next token, we prune the retrieval tokens with a simple heuristic strategy: if the rank of the next next token is below than 8, we preserve 4 tokens; if the rank of the next next token is below than 32, we preserve 3 tokens; otherwise we only preserve the speculate next next token itself. The retrieval process is terminated until the total number of draft tokens exceeds a specific draft tree capacity ${K}$.

### v2

@@ Abstract
[…]
LogitSpec: Accelerating Retrieval-based Speculative Decoding via Next Next Token Speculation
Tianyu Liu1,2thanks: Equal Contribution.   Qitan Lv1,2∗   Hao Li1,2   Xing Gao2   Xiao Sun2thanks: The Corresponding Author.   Xiaoyan Sun1 1University of Science and Technology of China  2Shanghai AI Laboratory {tianyu_liu, qitanlv, lihaohn}@mail.ustc.edu.cn sunxiao@pjlab.org.cnsunxiaoyan@ustc.edu.cn
@@ 1 Introduction
To address this challenge, speculative decoding (SD) (Leviathan et al., 2023; Chen et al., 2023) has emerged as a promising approach for lossless LLM inference acceleration. The key idea of SD is to employ a small draft model to first generate $\gamma$ draft tokens, and then the target model validates these tokens in parallel within a single forward, ensuring the output distribution to be unchanged. However, deploying an extra draft model introduces considerable overhead and difficulties, including (a) complex implementation and deployment, especially in integration with other techniques,(Liu et al., 2025) (b) additional memory overhead, especially in long-context scenarios. Moreover, when there exists no available draft model for SD, it takes substantial training cost to construct a compact draft model (Li et al., 2025b). (Please refer to Section 2 for more details.)
Therefore, many existing works are dedicated to developing draft-model-free SD methods (Saxena, 2023; Fu et al., 2024; He et al., 2024), where the draft tokens are generated in a retrieval-based manner. At each decoding step, they retrieve the relevant reference according to the last few tokens and extract draft tokens from the reference. In this way, they can explicitly eliminate the need for a draft model and reduce the drafting overhead.
Figure 1: Illustration of vanilla retrieval-based SD method and our LogitSpec. (a) Retrieval-based SD retrieves the wrong token “triangle” according to the next token “the”. (b) LogitSpec first speculates the next next token “area”, and then retrieves the right relevant reference “area of” according to “the area”. This simple example illustrates how LogitSpec utilizes the last logit to speculate the next next token and improves the retrieval accuracy.
@@
In this paper, we propose a simple yet effective retrieval-based SD framework, namely LogitSpec, which leverages the logit of the last token (last logit) to predict the next next token and improves the accuracy of the retrieval reference. LogitSpec is motivated by the observation that the last logit can speculate the next next token with a relatively high accuracy. Based on this observation, we use the speculated next next token as guidance to retrieve reference. Specifically, LogitSpec generates draft tokens in two steps: (1) utilizing the last logit to speculate the next next token; (2) retrieving relevant reference for both the next token and the next next token. As shown in Figure 1 (b), with the speculated token “area”, LogitSpec successfully retrieves the correct draft tokens “area of”. This two-step process enables better retrieval accuracy—it can help filter the most relevant reference when the reference is redundant, while extending the searching space when there exists no relevant reference. To summarize, our contributions are:
@@
    We empirically observe that the last logit can speculate the next next token with a relatively high accuracy. Unlike other retrieval techniques, which are sensitive to the specific task, this property is robust and effective across various tasks.
@@
[…]
@@ 2 Related Work
This line of research focuses on improving the draft quality, which primarily invests heavily in post-training a specialized draft model to generate high-quality drafts. Medusa (Cai et al., 2024) pioneers this direction by adding extra MLP heads at the top of the target model, reusing the hidden states from the target model, and predicting the next few tokens in parallel. However, the generation from these decoding heads is independent, which harms the accuracy of the draft tokens. Therefore, some works (Ankner et al., 2024; Xiao et al., 2024) propose to add sequential dependencies for better performance. Glide (Du et al., 2024) takes a similar idea of Medusa and proposes to reuse the KV cache from the target model. Recently, Eagle (Li et al., 2025c, 2024, b) has dominated this research line by training an auto-regressive head and scaling up the training data. However, while these methods achieve superior speedup, all of them rely on an extra draft model, which necessitates extra parameters or extensive training. The existence of the draft model significantly limits its application: (a) deploying a draft model requires complex and careful implementation. (b) Both the draft model and its KV cache require additional GPU memory, which may incur load imbalance. In the long context settings, even the KV cache for the draft model may exceed the capacity of 1 GPU and lead to significant resource competitions; (c) the training cost is substantial, and the draft model requires retraining when the target model is updated.
@@
This line of research focuses on maximizing drafting efficiency by eliminating the need for a draft model entirely, aiming for universal, zero-cost acceleration. Some existing works (Zhang et al., 2024a; Elhoushi et al., 2024; Xia et al., 2025) observe that the layer sparsity of LLMs means that not all the layers are necessary to predict the next token. They propose methods to generate draft tokens with a subset of the layers of the target model. However, the number of skipped layers is limited, which affects the overall speedup. Recently, retrieval-based SD (He et al., 2024; Saxena, 2023; Fu et al., 2024) has dominated the draft-model-free speculative decoding by retrieving the relevant reference as the draft tokens. The relevant reference can be derived from the question and generated tokens, or an external database. However, there exist two common issues during the retrieval process: (a) retrieval from the context often fails to find matched draft tokens; (b) retrieval from an external database often struggles to find accurate draft tokens. These two issues motivate us to develop a more appropriate retrieval-based speculative decoding framework.
@@ 3 Background
Let $\bm{x}$ denote an input sequence (prefix). A speculative decoding step consists of a drafting phase and a verification phase. During the drafting phase, the draft model $\mathcal{M}_{q}$ is employed to give $\gamma$ draft tokens $x_{1},x_{2},...,x_{\gamma}$ by running $\gamma$ times the model forward and sampling. Here, we denote the output logit $\mathcal{M}_{q}(\bm{x}+[x_{1},...,x_{i-1}])$ as $q_{i-1}$, then each draft token is given by $x_{i}\sim q_{i-1},i=1,...,\gamma$. During the verification phase, the prefix $\bm{x}$ together with $\gamma$ draft tokens are sent to $\mathcal{M}_{p}$ for verification. The target model $\mathcal{M}_{p}$ inputs $\bm{x}+[x_{1},...,x_{\gamma}]$ and outputs the logits $p_{0},p_{1},...,p_{\gamma}$. Then SD sequentially verifies $x_{i}$ via speculative sampling (Leviathan et al., 2023; Chen et al., 2023), where the acceptance rate is given by:
[…]
@@
$\mathcal{R}=\{(x_{i}^{1},x_{i}^{2},...,x_{i}^{n})\}_{i=1}^{N}.$ (2)
Here, $N$ denotes the total number of $n$-grams in $\mathcal{R}$. Existing retrieval-based SD methods mainly differ in constructing $\mathcal{R}$. At each SD decoding step, given a query $m$-gram tokens $x_{q}^{1},x_{q}^{2},...,x_{q}^{m}$, the retrieval model first traverses the reference, finds matched $n$-grams and returns the subsequent tokens of the matched tokens:
$\textsc{Match}\big(\mathcal{R},(x_{q}^{1},x_{q}^{2},...,x_{q}^{m})\big)\\ =\{(x_{i}^{m+1},...,x_{i}^{n})\mid i\in\mathcal{S}\}.$ (3)
@@ 4 Method
Figure 2: Motivated observation I: the last logit can speculate the next next token with high accuracy. In over 50% decoding steps, the next next token can be found in the top-60 entries within the last logit across model sizes and architectures.
@@ 4.1 Motivated Observation
Drawing insights from multi-token prediction (Gloeckle et al., 2024) that fine-tuned LLMs can predict multiple tokens in a single forward, we are interested in the ability of LLMs to predict the next next token without fine-tuning. We conduct a simple experiment to investigate the rank of the next next token in the last logit, which is used to predict the next token. Specifically, we use a small fraction of Spec-Bench (Xia et al., 2024) that contains 13 sub-tasks (detailed in Appendix B) and 6 different LLMs as backbones. Results in Figure 2 demonstrate that the last logit has a strong potential to predict the next next token. In over 50% decoding steps, the next next token can be found within the top-60 entries of the last logit. For Llama 3.1 8B (Team, 2024) with a large 128K vocabulary, this percentage even increases to 64% within the top-60 entries of the last logit. The results of 6 different LLMs on 13 downstream tasks demonstrate that this observation holds consistently across different models (Vicuna, Llama 3.1, and Qwen) and scales (ranging from 0.5B to 33B).
Figure 3: Motivated observation II: compared with other retrieval-based methods, the prediction of the last logit demonstrates robustness to downstream tasks, motivating us to utilize it to guide the retrieval process.
We further conduct experiments on Spec-Bench with Vicuna 7B to investigate the robustness of the predictive capability of the last logit across different tasks. We propose a minimal speculative decoding algorithm with last logit, last logit decoding for this experiment, where we directly use the top-60 entries of the last logit to serve as the draft tokens. Each draft token is a guess at the next next token. If the target model accepts any one of the draft tokens, the others are dropped. Consequently, at each decoding step, the accepted length is either 2 or 1. As in Figure 3, this simple SD method yields mean accepted tokens per decoding step (MAT) of over 1.5. Compared with other retrieval-based SD methods such as PLD (Saxena, 2023) and REST (He et al., 2024), which are highly affected by the task type, the MAT of last logit decoding exhibits superior robustness to the task. These results demonstrate that the ability of the last logit to predict the next next token widely exists in modern LLMs themselves, and strongly motivate us to improve retrieval-based SD methods with the last logit. We provide a detailed experimental setup for the motivation experiments in Appendix B.
Figure 4: An overview of LogitSpec. At each decoding step, LogitSpec first utilizes the top-$k$ entries of the last logit as the speculation for the next next token. Then, LogitSpec retrieves relevant references for both the next token and the next next token. Finally, LogitSpec organizes the draft tokens into a draft tree and prepares a tree attention for parallel verification.
@@ 4.2 LogitSpec Drafting
While last logit decoding has shown great simplicity in implementation and robustness to the downstream tasks, its theoretical upper bound of the speedup ratio is 2, as all the draft tokens are guesses for the next next token. In our experiments, the last logit decoding achieves an overall speedup ratio of $1.2\times\sim 1.4\times$, which hinders the real-world applications of last logit decoding due to the unsatisfactory speedup ratio. Therefore, to further improve the overall inference speedup, we propose to utilize the prediction ability of the last logit as a guidance for retrieval. Specifically, for a given input prefix $\mathbf{x}=(x_{1},x_{2},...,x_{i})$ and the last logit $p_{i}=\mathcal{M}(x_{1},x_{2},...,x_{i})$, we first sample the next token $x_{i+1}\sim p_{i}$. Then, we utilize the top-$k$ entries of the last logit (except the next token $x_{i+1}$) as the speculation to the next next token $\widetilde{x}_{i+2}$:
$\mathcal{L}=\{\widetilde{x}_{i+2}\mid\widetilde{x}_{i+2}\in\textsc{Top}_{k}(p_{i}),\,\widetilde{x}_{i+2}\neq x_{i+1}\}.$ (4)
Then, we use the retrieval model $\mathcal{R}$ to retrieve reference for both the next token and next next tokens:
$\mathcal{D}_{i+1}(x_{i+1})=\textsc{Match}\big(\mathcal{R},\bm{g}^{(1)}_{i}(x_{i+1})\big),$ (5)
$\mathcal{D}_{i+2}(\widetilde{x}_{i+2})=\textsc{Match}\big(\mathcal{R},\bm{g}^{(2)}_{i}(\widetilde{x}_{i+2})\big),$ (6)
$\mathcal{D}=\mathcal{D}_{i+1}(x_{i+1})\cup\bigcup_{\widetilde{x}_{i+2}\in\mathcal{L}}\mathcal{D}_{i+2}(\widetilde{x}_{i+2}).$ (7)
Here, $\bm{g}^{(1)}_{i}(x_{i+1})=(x_{i-m+2},...,x_{i},x_{i+1})$ denotes the $m$-gram query for the next token, and $\bm{g}^{(2)}_{i}(\widetilde{x}_{i+2})=(x_{i-m+3},...,x_{i},x_{i+1},\widetilde{x}_{i+2})$ denotes the query for the next next token. In our experiments, we retrieve draft tokens with $m=3$ first. If no matched tokens are found, we decrease $m$ to 2 and so on. For simplicity, we only use the user-input prompt and decoded tokens to construct the retrieval model $\mathcal{R}$, that is to say, our LogitSpec is query-independent, and will not be affected by other queries. We give the detailed implementation of the retrieval process in Appendix D.3.
Taking the example in Figure 4, we first select the top-$k$ entries of the last logit as the speculation for the next next token. Then, we retrieve reference for both the next token “the” and the next next token “area”, and return “of the”. After that, we organize these draft tokens into a draft tree, where the next token serves as the root. Finally, the token sequence “area of the” is accepted by $\mathcal{M}$.
After the retrieval process, we employ a simple pruning strategy to control the number of draft tokens. Specifically, we do not prune the retrieval tokens for the next token. For each speculated next next token, we prune the retrieval tokens with a simple heuristic strategy: if the rank of the next next token is below 8, we preserve 4 tokens; if the rank of the next next token is below 32, we preserve 3 tokens; otherwise, we only preserve the speculated next next token itself. The retrieval process is terminated until the total number of draft tokens exceeds a specific draft tree capacity K{K}. Furthermore, we explored the impact of various pruning strategies, with results detailed in Table 14 of Appendix E.7.